OpenStudy (this):

some on help me plllz a sphere has a surface area of 32in^2 a,radius=_______ b,volume=_________

OpenStudy (ajprincess):

Surface area of a sphere = \(4\pi r^2\) \(2=4\pi r^2\) From this r(radius) can be found in terms of \(\pi\) Volume of a sphere=\(\large\frac{4}{3}\pi r^3\) Can u find nw? @this

OpenStudy (this):

sa=4pi r^2 32/4pi=4pir^2/4pi 32=r^2 i dont know what to do next):

OpenStudy (ajprincess):

32/4pi=4pi r^2/4pi 8/pi=r^2 to find r. find the sqrt of 8/pi r=sqrt(8/pi)

OpenStudy (ajprincess):

to find the volume use v=4/3pi r^3

OpenStudy (this):

thanks so much @ajprincess

OpenStudy (ajprincess):

Welcome:) Did u find the volume?

OpenStudy (this):

i got 2.25=r is that right?

OpenStudy (ajprincess):

pi=?

OpenStudy (this):

i cancel out the pi

OpenStudy (ajprincess):

what did use for pi? 3.14 or 22/7?

OpenStudy (this):

i use 3.48

OpenStudy (this):

wait let me show u how i did it the whole thing

OpenStudy (ajprincess):

ok:)